OpenStudy (anonymous):

b/3 + 1/8 = 19 step by step thanks

OpenStudy (anonymous):

minus both side by 1/8 u'll have b/3=151/8 now cross multiply

OpenStudy (anonymous):

by cross multiply we have 8(b)=151(3)

OpenStudy (anonymous):

then solve for b

OpenStudy (anonymous):

8b=483 find b